# A theory of ecological invasions and its implications for eco-evolutionary dynamics

### v1

@@ I Introduction
Despite the importance of ecological invasions, our theoretical understanding remains limited [13, 14, 15, 16]. This gap in our knowledge is even more pronounced when the communities of interest are diverse, evolved, and/or engage in complex behaviors such as cross-feeding [17]. Invasion outcomes are mediated by a variety of factors including invader characteristics, community interactions, and environmental conditions [18, 5, 19, 20]. A further complication that arises when thinking about invasions is that invaders fundamentally alter the ecosystem they invade [21, 22].
@@ II Theory of ecological invasions
In this work, we introduce a general framework to predict the new ecological steady state after an invasion, making use of the knowledge of the pre-invasion community and their interactions with the invader. The outcomes of successful invasions can be categorized by three experimentally measurable quantities: (1) the identity of species that go extinct, (2) the shifts in abundance among surviving species, and (3) the abundance of the invader itself. When the pre-invasion ecosystem is large and stable, invasions can be viewed as small perturbations to the existing ecosystems. Our framework exploits this observation to predict these outcomes systematically by developing a generalized linear-response theory for ecological invasions. Even though our assumption of invaders being small perturbations is justified only for large communities, our framework nevertheless successfully predicts invasion outcomes for small communities as well (see section V).
@@ II.1 Predicting invasion outcomes
where $i$ labels species from 1 to $S$, $X_{i}$ is the abundance of species $i$ and $g_{i}(\vec{X})$ is its net per capita growth rate in the presence of other species. $g_{i}(\vec{X})$ is a linear function of its naive growth rate $r_{i}$ and the abundances of all species $X_{j}$. The interaction coefficients $A_{ij}$ characterize how the presence of species $j$ affects the growth rate of species $i$.
Throughout this study we assume that the dynamics converge to a stable steady state. Denote the pre-invasion abundances of the species in the community by ${X}^{\text{\text{old}}}_{j}$. We then ask what happens when we invade such an ecosystem by a new species, denoted by the distinguished index 00. Within GLV, the invader is fully characterized by specifying its interaction coefficients $A_{0j}$ and $A_{j0}$ and its growth rate $r_{0}$.
@@
To achieve our goal, we will write a set of perturbative self-consistency equations for our quantities of interest. Assume we know the outcome of the invasion, and in particular which species survive, $\mathcal{S}$, and which species go extinct, $\mathcal{E}$. Then we can split the pre-invasion interaction matrix of the community into four blocks,
@@
Our key insight is to note that even though the abundances of extinct species $X_{\mathcal{E}}$ change abruptly to zero post-invasion, the post-invasion growth rates of extinct species $g_{\mathcal{E}}(\vec{X}^{\text{new}})$ change more smoothly from positive to negative and can hence still be calculated perturbatively. A related observation undergirds the use of slack variables in constrained optimization problems with inequality constraints [33]. Drawing on these observations, in Appendix A.3 we show the invader abundance can be expressed in terms of the quantities introduced above as
@@
To close this set of equations, we impose self-consistency. Namely, we require that surviving species $\mathcal{S}$ have positive abundances
@@
Collectively, Eqs. (4)-(7) define a set of self-consistent equations (a mix of equalities and inequalities) that must be satisfied by the post-invasion abundances. These equations play a role similar to the optimality equations in constrained optimization and hence can be solved iteratively by introducing auxiliary slack variables. Details of this procedure are outlined in the Appendix. We also show below that these equations have an natural ecological interpretation in terms of a new ecological quantity we call the “dressed invasion fitness” that generalizes the notion of invasion fitness to account for community-mediated ecological feedbacks.
@@ II.2 Tests of model predictions
Figure 2: Theory makes accurate predictions across diverse ecological models. Comparison between simulations theoretical predictions for (a–c) Lotka-Volterra models, (d–g) MacArthur consumer resource models, and (h–k) microbial consumer resource models with cross-feeding. Panels display simulated versus predicted abundances for invaders, community residents and resources respectively. Each data points indicate one species/resource. Insets for each panel show (1) matrices indicating qualitative prediction outcomes (extinct or surviving; rows indicate simulations, columns predictions); and (2) histograms of relative errors.
In the last section, we derived a set of self-consistency equations describing ecological invasions for the Lotka-Volterra model. In Appendix E, we show that self-consistency equations analogous to Eqs. (4)-(7) can be derived for a large class of niche-based ecological models, including most variants of consumer resource models. In this more general setting, the pairwise interaction matrix $A$ is replaced by an inverse of a susceptibility matrix that captures how a perturbation in the abundance of one species affects the abundances of other species. In addition, for consumer resource models, the self-consistency equations for species must be augmented by additional equations describing the response of resources to invasion. Nonetheless, the qualitative logic, derivation, and structure of the self-consistency equations remains the same.
To demonstrate the generality of our framework, we compare our theoretical predictions to numerical simulations for three different classes of niche-based models in theoretical ecology: (i) the Generalized Lotka-Volterra model, (ii) the MacArthur consumer resource model ; and the (iii) Microbial consumer resource model with cross-feeding (see Fig. 2).
Lotka-Volterra models. We simulated $640$ communities with $10$ species and dynamics given by the generalized Lotka-Volterra equations in Eq. (1). We focused on ecosystems with symmetric interactions where the interaction coefficients $A_{ij}$ were drawn from a random normal distribution (see Appendix H.1 for details of parameters). After running the dynamics until steady state, on average $8$ of the $10$ original species survived in our simulations. We then introduced an invader into the ecosystem and compared theoretical predictions to numerical simulations. We focused on three quantities: (a) predicting whether the invader would invade successfully, (b) the post-invasion invader abundance for successful invasions, and (c) the resulting shift in the species abundances (the community shift). As can be seen in Fig. 2 (b)-(c), the theory and simulations agree remarkably well.
@@
$\frac{dN_{i}}{dt}$ $=N_{i}\left(\sum_{\alpha}C_{i\alpha}R_{\alpha}-m_{i}\right)$ (8)
@@
To test these theoretical predictions, we simulated $640$ communities using Eq. (9), with parameters drawn randomly (see Appendix H.2 for details). We initialized each community with $60$ consumers and 30 resources and ran the dynamics to steady state. After this initial phase, we introduced an invader and simulated the post-invasion ecosystem until it reached a steady state. We then compared our theoretical predictions for invasion success, invader abundance, and the shifts in species and resource abundances with the results of our simulation (see Fig. 2 (e)-(g)). As in the Lotka-Volterra model, our theory for the MCRM recapitulates the simulation results.
Microbial consumer resource model with cross-feeding. Finally, to test our theory in a more complex ecological setting, we applied our framework to the Microbial consumer resource model (MiCRM) [31, 36]. The MiCRM extends traditional consumer resource models by explicitly modeling cross-feeding by microbial species — where metabolic byproducts produced by one species are consumed by others. As a result of this cross-feeding, the steady-state conditions of the MiCRM are non-linear. We show in the Appendix that we can still perturbatively derive self-consistency equations for invasion success, invader abundance, and the shifts in species and resource abundances. A detailed discussion of the MiCRM and this calculation can be found in Appendix E. Fig. 2(i-k) shows a comparison with good agreement between theory and simulation (see Appendix H.3 for numerical detail). These results suggest that our theory applies to a variety of ecological models and successfully predicts all key properties of invasion outcomes.
@@ III Dressed invasion fitness governs outcomes
Figure 3: Dressed invasion fitness and extinction-induced feedbacks govern invasion outcomes. Our theory shows that both (a) invader abundance and (b) shifts in community abundances can be decomposed into two parts: a direct contribution the invader — through the dressed invasion fitness — and an indirect contribution due to extinctions (Eqs (11)). (c–i) Both Lotka-Volterra and Microbial consumer resource model simulations that (c–d) invader abundance is generally correlated with naive invasion fitness, but (d) and (g) show that dressed invasion fitness — which correctly incorporates ecological feedbacks — is a much better predictor of invader abundance. Similarly (e) and (h–i) show that our theory can also generally predict shifts in community abundances through the invader’s direct effect (dashed line and blue dots) in the absence of extinctions. To predict the effect of invaders that introduce more extinctions — which show deviations from direct effect-based predictions (colored dots) — we need to account for indirect effects as well (Eq. (6)).
@@
Consider an invading an ecosystem governed by Eq. (1) by a new species 00 with naive growth rate $r_{0}$ and interaction matrix $A_{0j}$. The invasion fitness $g_{0}^{\text{naive}}$ of species 00 is given by
@@
Eq. (11) shows that the invader abundance post-invasion can be decomposed into a sum of two terms: (1) a term proportional to the invasion fitness, which we call the dressed invasion fitness $g^{\text{dressed}}_{0}$; and (2) a term due to species extinctions $\vec{c}_{\mathcal{E}}\cdot\vec{X}_{\mathcal{E}}^{\text{old}}$ (Fig.  3a). The dressed fitness, unlike naive fitness $g_{0}^{\text{naive}}$, accounts for the propagation of feedback from invader to surviving community and back to invader (the second term in the denominator of Eq. (12)). Figs. 3 (c-d) show the invader abundance $X_{0}$ as a function of invasion fitness $g_{0}^{\text{naive}}$ and dressed invasion fitness $g_{0}^{\text{dressed}}$. As expected, both these quantities are correlated with invader abundance. The dressed invasion fitness is much more predictive of the final invader abundance, especially in the absence of species extinctions (Fig. 3d). This shows that the dressed invasion fitness which incorporates ecological feedbacks is really the “right” quantity to predict the fate of an invader.
@@
Analogous interpretations exist for consumer-resource models, including the MiCRM (see Appendix D -F). Once again all relevant quantities can be decomposed into a contribution due to the invader and a contribution due to extinctions. As can be seen in Figs. 3 (f–i), these decompositions continue to hold in consumer resource models despite the presence of more complex dynamics.
@@ IV Implications for eco-evolutionary dynamics
Our theoretical framework is especially well-suited for analyzing short-term evolutionary dynamics of this type because it makes no assumptions about invader characteristics or the ecosystem being invaded. This flexibility allows us to use our theoretical results to model eco-evolutionary processes despite the fact that evolution often generates complex correlations between species that co-exist in the ecosystem. An example of this is shown in Fig. 4(b) which shows how the consumer preferences of species in the MiCRM change as a community is evolved by successive attempted invasions by highly-correlated mutants (Methods). Remarkably, after just $40$ successful mutations, the consumer preference matrices of the evolved and assembled communities looked qualitatively distinct (Fig, 4(b)). Quantifying these differences revealed two key distinctions: first, species in evolved communities exhibit much higher trait correlations, with the inter-species correlation histogram showing a second peak near 0.8 (Fig. 4(e)); second, evolved communities contain fewer low-abundance species (Fig. 4(d)). This second pattern emerges because low-abundance species disproportionately go extinct following successful mutations (Fig. 4(c)), regardless of whether communities are evolved or assembled.
Given the very different structure of evolved and assembled communities and the fact that mutants are highly correlated with species that already exist in the community, we wanted to directly test whether our theoretical framework could be used to predict the outcome of a single step of evolution. To do so, we simulated the evolution of 640 distinct communities and then introduced a new mutant into each of these evolved communities (see Appendix H.4). We then used our theory to predict whether the the mutant would fix, and if it fixed, the abundance of the mutant as well as the abundances of the species and resources in the community. Despite the complexities introduced by evolutionary history, Fig. 4(f-h) shows that our theoretical predictions show remarkable agreement with simulations. These results demonstrate that evolutionary dynamics can be captured — despite complex evolutionary histories, mutant-parent correlations, and community diversity — as long as one incorporates ecological feedbacks.

### v2

@@ I Introduction
Despite the importance of ecological invasions, our theoretical understanding remains limited [13, 14, 15, 16]. This gap in our knowledge is even more pronounced when the communities of interest are diverse, evolved, and/or engage in complex behaviors such as cross-feeding [17]. Invasion outcomes are mediated by a variety of factors including invader characteristics, community interactions, and environmental conditions [18, 5, 19, 20]. A further complication that arises when thinking about invasions is the presence of ecological feedbacks, i.e., the fact that invaders fundamentally alter the ecosystem they invade [21, 22].
@@ II Theory of ecological invasions
In this work, we introduce a general framework to predict the new ecological steady state after an invasion, making use of the knowledge of the pre-invasion community and their interactions with the invader. The outcomes of successful invasions can be categorized by three experimentally measurable quantities: (1) the identity of species that go extinct, (2) the shifts in abundance among surviving species, and (3) the abundance of the invader itself. When the pre-invasion ecosystem is sufficiently diverse (see Fig.S4) and at steady-state, the introduction of invaders can be viewed as adding a small initial perturbation to the existing ecosystem. Our framework exploits this observation to predict invasion outcomes systematically using a linear-response theory for ecological invasions, which works even though invaders that start at low abundances can eventually grow to large abundances. Our assumption of linear response is justified whenever the effective ecological interactions in pre- and post-invasion communities are approximately similar (see Appendix C for precise definition). For many canonical ecological models (e.g., the Lotka-Volterra and MacArthur consumer resource models), we can show that this assumption is exactly true, not just approximate (Appendix C and E). For a wide variety of other models, we find that this assumption is often approximately valid (see Fig. 2 and Fig. S4). As a result, our framework can successfully predict invasion outcomes even in experimental data from low-diversity communities with $10\%$ error (see Section V).
@@ II.1 Predicting invasion outcomes
where $i$ labels species from 1 to $S$, $X_{i}$ is the abundance of species $i$ and $g_{i}(\vec{X})$ is its net per capita growth rate in the presence of other species. $g_{i}(\vec{X})$ is a linear function of its naive growth rate $r_{i}$ and the abundances of all species $X_{j}$. The interaction coefficients $A_{ij}$ characterize how the presence of species $j$ affects the growth rate of species $i$. Throughout this study we assume that the dynamics converge to a steady state, but it need not be unique. Our framework works equally well for ecosystems with multiple stable steady states (see Fig. S2). Denote the pre-invasion abundances of the species in the community by ${X}^{\text{\text{old}}}_{j}$. We then ask what happens when we invade such an ecosystem by a new species, denoted by the distinguished index 0. Within GLV, the invader is fully characterized by specifying its interaction coefficients $A_{0j}$ and $A_{j0}$ and its growth rate $r_{0}$.
@@
To achieve our goal, we will write a set of self-consistency equations for our quantities of interest. Assume we know the outcome of the invasion, and in particular which species survive, $\mathcal{S}$, and which species go extinct, $\mathcal{E}$. Then we can split the pre-invasion interaction matrix of the community into four blocks,
@@
Our key insight is to note that even though the abundances of extinct species $X_{\mathcal{E}}$ change abruptly to zero post-invasion, the post-invasion growth rates of extinct species $g_{\mathcal{E}}(\vec{X}^{\text{new}})$ change more smoothly from positive to negative and can hence still be calculated perturbatively.
Drawing on these observations, in Appendix A.3 we show the invader abundance can be expressed in terms of the quantities introduced above as
@@
Note that we do not assume that invaders always cause a weak effect on a community: we explicitly include the strong effects of invaders on communities over the invasion process. Invaders exert a weak effect on the community initially, i.e., they invade at low abundance, as is common assumption across ecological literature [33]. However, as invaders grow their effect becomes strong and can cause large changes to communities, such as extinctions of other resident species. These effects appear explicitly in Eqs. (5)–(6), as surviving species $\mathcal{S}$ behave differently from extinct ones $\mathcal{E}$. To close this set of equations, we impose self-consistency. Namely, we require that surviving species $\mathcal{S}$ have positive abundances
@@
Collectively, Eqs. (4)-(7) define a set of self-consistent equations (a mix of equalities and inequalities) that must be satisfied by the post-invasion abundances. Fully solving these equations requires an exhaustive combinatorial scan that is computationally expensive. Instead, we use a simpler iterative approach (details in Appendix I Algorithm. 1) which is common in constrained optimization because of its efficiency and accuracy ($\sim 99\%$ in our case). We will later show that these self-consistency equations have an natural ecological interpretation in terms of a new ecological quantity we call the “dressed invasion fitness” that generalizes the notion of invasion fitness to account for community-mediated ecological feedbacks.
@@ II.2 Tests of model predictions
Figure 2: Theory makes accurate predictions across diverse ecological models. Comparison between simulations theoretical predictions for (a–c) Lotka-Volterra models, (d–g) MacArthur consumer resource models with Type I functional response, (h–k) Monod consumer resource with Type II functional response (Monod dependence of growth rate on resource abundances), and (l–o) microbial consumer resource models with cross-feeding. Panels display simulated versus predicted abundances for invaders, community residents and resources respectively. Each data point indicates one species/resource. Insets for each panel show (1) matrices indicating qualitative prediction outcomes (extinct or surviving; rows indicate simulations, columns predictions); and (2) histograms of relative errors.
In the last section, we derived a set of self-consistency equations describing ecological invasions for the Lotka-Volterra model. In Appendix F, we show that self-consistency equations analogous to Eqs. (4)-(7) can be derived for a large class of niche-based ecological models, including most variants of consumer resource models. In this more general setting, the pairwise interaction matrix $A$ is replaced by an inverse of an effective interaction matrix that captures how a perturbation in the abundance of one species affects the abundances of other species. In addition, for consumer resource models, the self-consistency equations for species must be augmented by additional equations describing the response of resources to invasion. Nonetheless, the qualitative logic, derivation, and structure of the self-consistency equations remains the same.
To demonstrate the generality of our framework, we compare our theoretical predictions to numerical simulations for four different classes of niche-based models in theoretical ecology: (i) the Generalized Lotka-Volterra model, (ii) the MacArthur consumer resource model with type I functional response; the (iii) the Monod consumer resource model with type II functional response, and (iv) the Microbial consumer resource model with cross-feeding (see Fig. 2).
Lotka-Volterra models. We simulated $640$ communities with $10$ species and dynamics given by the generalized Lotka-Volterra equations in Eq. (1). We focused on ecosystems with symmetric interactions where the interaction coefficients $A_{ij}$ were drawn from a random normal distribution (see Appendix K.1 for details of parameters). After running the dynamics until steady state, on average $8$ of the $10$ original species survived in our simulations. We then introduced an invader into the ecosystem and compared theoretical predictions to numerical simulations. We focused on three quantities: (a) predicting whether the invader would invade successfully, (b) the post-invasion invader abundance for successful invasions, and (c) the resulting shift in the species abundances (the community shift). As can be seen in Fig. 2 (b)-(c), the theory and simulations agree remarkably well.
@@
$\frac{dN_{i}}{dt}$ $=N_{i}\left(\sum_{\alpha}C_{i\alpha}R_{\alpha}-m_{i}\right)$
@@
To test these theoretical predictions, we simulated $640$ communities using Eq. (8), with parameters drawn randomly (see Appendix K.2 for details). We initialized each community with $60$ consumers and 30 resources and ran the dynamics to steady state. After this initial phase, we introduced an invader and simulated the post-invasion ecosystem until it reached a steady state. We then compared our theoretical predictions for invasion success, invader abundance, and the shifts in species and resource abundances with the results of our simulation (see Fig. 2 (e)-(g)). As in the Lotka-Volterra model, our theory for the MCRM recapitulates the simulation results. The error distributions for this model (insets of Figs. 2(e–g)) have two modes: the left peak ($\sim 10^{-6}\%$) is due to numerical errors, while the right peak ($\sim 1\%$) is due to occasional errors in our iterative approach to solve the self-consistency equations (a non-iterative approach removes it; see Fig. S6).
Monod model and Microbial consumer resource model with cross-feeding. Finally, to test our theory in more complex ecological settings, we applied our framework to two additional models: a consumer resource model with Monod or type II functional response, and the Microbial consumer resource model (MiCRM) [31, 36]. The Monod model extends the MacArthur model to include a nonlinear dependence of species’ per capita growth rates on resource abundances (Appendix F). The MiCRM extends traditional consumer resource models by explicitly modeling cross-feeding by microbial species — where metabolic byproducts produced by one species are consumed by others. As a result of these extensions, the effective interaction matrix $A^{\text{eff}}$ in these models depend on community state and are not fixed. In these cases, our theory assumes that the post-invasion $A^{\text{eff}}$ is still approximately the same as the one pre-invasion. We can thus use the same steps as before to derive self-consistency equations for invader abundance and community abundance shifts (details of approximations and derivation for both models in Appendix F). Fig. 2(i–o) shows that despite approximations, there is remarkably good agreement between theoretical predictions and simulations (simulation details in Appendix K.3). These results suggest that our theory applies to a variety of ecological models and successfully predicts all key properties of invasion outcomes.
@@ III Dressed invasion fitness governs outcomes
Figure 3: Dressed invasion fitness and extinction-induced feedbacks govern invasion outcomes. Our theory shows that both (a) invader abundance and (b) shifts in community abundances can be decomposed into two parts: a direct contribution the invader — through the dressed invasion fitness — and an indirect contribution due to extinctions (Eqs (11)). (c–i) Both Lotka-Volterra and Microbial consumer resource model simulations that (c–d) invader abundance is generally correlated with naive invasion fitness, but (d) and (g) show that dressed invasion fitness — which correctly incorporates ecological feedbacks — is a much better predictor of invader abundance. Similarly (e) and (h–i) show that our theory can also generally predict shifts in community abundances through the invader’s direct effect (dashed line and blue dots) in the absence of extinctions. To predict the effect of invaders that introduce more extinctions — which show deviations from direct effect-based predictions (colored dots) — we need to account for indirect effects as well (Eq. (6)). Fig. S3 shows analogous plots for the the MacArthur and Monod consumer resource models.
@@
Consider an invading an ecosystem governed by Eq. (1) by a new species 0 with naive growth rate $r_{0}$ and interaction matrix $A_{0j}$. The invasion fitness $g_{0}^{\text{naive}}$ of species 0 is given by
@@
Eq. (10) shows that the invader abundance post-invasion can be decomposed into a sum of two terms: (1) a term proportional to the invasion fitness, which we call the dressed invasion fitness $g^{\text{dressed}}_{0}$; and (2) a term due to species extinctions $\vec{c}_{\mathcal{E}}\cdot\vec{X}_{\mathcal{E}}^{\text{old}}$ (Fig.  3a). The dressed fitness, unlike naive fitness $g_{0}^{\text{naive}}$, accounts for the propagation of feedback from invader to surviving community and back to invader (the second term in the denominator of Eq. (11)). While the naive fitness only measures the invader’s initial growth rate at small abundance, the dressed invasion fitness accounts for the invader’s effect on the community as its abundance gets larger. As an invader grows, it starts to affect other community members, which affect it back. The dressed invasion fitness augments naive invasion fitness (numerator) with exactly this effect (denominator of Eq. (11)). Figs. 3 (c-d) show the invader abundance $X_{0}$ as a function of invasion fitness $g_{0}^{\text{naive}}$ and dressed invasion fitness $g_{0}^{\text{dressed}}$. As expected, both these quantities are correlated with invader abundance. The dressed invasion fitness is much more predictive of the final invader abundance, especially in the absence of species extinctions (Fig. 3d). This shows that the dressed invasion fitness which incorporates ecological feedbacks is really the “right” quantity to predict the fate of an invader.
@@
Analogous interpretations exist for consumer-resource models, including the Monod model and MCRM (see Appendix E -H and Figs. S3, omitted here for brevity). Once again all relevant quantities can be decomposed into a contribution due to the invader and a contribution due to extinctions. As can be seen in Figs. 3 (f–i), these decompositions continue to hold in consumer resource models despite the presence of more complex dynamics.
@@ IV Implications for eco-evolutionary dynamics
Our theoretical framework is especially well-suited for analyzing short-term evolutionary dynamics of this type because it makes no assumptions about invader characteristics or the ecosystem being invaded. This flexibility allows us to use our theoretical results to model eco-evolutionary processes despite the fact that evolution often generates complex correlations between species that co-exist in the ecosystem. An example of this is shown in Fig. 4(b) which shows how the consumer preferences of species in the MiCRM change as a community is evolved by successive attempted invasions by highly-correlated mutants (Methods). Remarkably, after just $40$ successful mutations, the consumer preference matrices of the evolved and assembled communities looked qualitatively distinct (Fig, 4(b)). Quantifying these differences revealed two key distinctions. First, species in evolved communities exhibit much higher trait correlations, with the inter-species correlation histogram showing a second peak near 0.8 (Fig. 4(e)). This is reflected in a generic block-like structure in evolved consumer-resource interactions (Fig. 4(b))[38]. Second, evolved communities contain fewer low-abundance species (Fig. 4(d)). This pattern emerges because low-abundance species disproportionately go extinct following successful mutations (Fig. 4(c)), regardless of whether communities are evolved or assembled.
Given the very different structure of evolved and assembled communities and the fact that mutants are highly correlated with species that already exist in the community, we wanted to directly test whether our theoretical framework could be used to predict the outcome of a single step of evolution. To do so, we simulated the evolution of 640 distinct communities and then introduced a new mutant into each of these evolved communities (see Appendix K.4). We then used our theory to predict whether the the mutant would fix, and if it fixed, the abundance of the mutant as well as the abundances of the species and resources in the community. Despite the complexities introduced by evolutionary history, Fig. 4(f-h) shows that our theoretical predictions show remarkable agreement with simulations. These results demonstrate that evolutionary dynamics can be captured — despite complex evolutionary histories, mutant-parent correlations, and community diversity — as long as one incorporates ecological feedbacks.